OpenStudy (anonymous):

The 3rd degree Taylor polynomial for cos x centered at a= pi/2 is given by cosx= -(x-(pi/2))+1/6 ((x-(pi/2))^3 +R3(x) Using this, estimate cos 87°correct to five decimal places.

OpenStudy (anonymous):

convert \[87^{\circ}\]to radians and substitute in for x and compute

OpenStudy (anonymous):

\[87^{\circ}\cdot \frac{ \pi }{ 180^{\circ} }=\frac{ 87 \pi }{ 180 }\]\[\left( \frac{ 87 \pi }{ 180 }-\frac{ \pi }{ 2 } \right)=\left( \frac{ 87 \pi }{ 180 }-\frac{90 \pi }{ 180 } \right)=-\frac{ 3 \pi }{ 180 }=-\frac{ \pi }{ 60 }\]

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

you're welcome